Lines

an addictive line game

Version 2.2 4 January 1994

by Mika Kortelainen

Lines is completely done by Mika Kortelainen and is copyrighted 1992–1994.

1 Copyrights and fees

Lines is completely done by Mika Kortelainen and is copyrighted 1992–1994. Lines is SHARE-WARE, suggested fee is $5 or 20–30 Finnish marks (FIM). Please, send only cash because cashing foreign cheques is more expensive than the suggested fee. Because this archive contains the full, un-restricted version of the game, do not wait any return mail from me. But, by indicating your interest, you will get me to continue to develop Lines further.

Lines uses several additional shared-libraries, which are copyrighted as follows:

               `medplayer.library' is \begin{matrix}c \\ ◯ \end{matrix} Teijo Kinnunen.
               `powerpacker.library' is \begin{matrix}c \\ ◯ \end{matrix} Nico François.

Amiga, Workbench and AmigaGuide are registered trademarks of Commodore-Amiga, Inc.

Other trademarks are the property of their owners.

2 Disclaimer

The author will not take any responsibility about caused harm or ruined mental health Lines may bring about. You use Lines at your own risk.

3 Distribution

Lines may not be distributed commercially without author’s written permission (be ready to pay something to get a permission). Lines may be distributed freely as long as no essential parts of the original archive are removed and no more than a nominal copying fee is required. Lines may be included in disk collections such as Fred Fish’s one and in CD-ROM collections, too.

The original archive consists of the following files:

‘Lines’ This is the game itself (PAL version).

‘LinesNTSC’

This is the NTSC version of the game.

‘Lines.info’ and ‘LinesNTSC.info’

These are the corresponding icons for easy startup from Workbench.

‘Lines.doc’ and ‘Lines.doc.info’

This file as an ANSI text file.

‘Lines.dvi’ and ‘Lines.dvi.info’

This file as a DVI-file.

‘Lines.guide’ and ‘Lines.guide.info’

This file as an AmigaGuide file. Uses features of AmigaGuide V39, but can be read using older AmigaGuide, too.

‘libs/medplayer.library’
‘libs/powerpacker.library’

The additional shared-libraries.

The NTSC version has some distortions in logos and such, but the actual playing board should be fine (I hope).

4 Requirements and installation

An Amiga computer (anything from 500 to 4000), TV or monitor, and a mouse are required to play Lines successfully. Lines uses ‘medplayer.library’ to play MED modules. ‘Powerpacker.library’ is needed if you intend to use packed modules. Most of you will already have these libraries in ‘LIBS:’ (at least owners of harddrives) but they are included in the archive and can be copied if needed. Besides, ‘printer.device’ is needed if you want to get hardcopies of your games (and also, you should have set the printer preferences using system software). Lines is tested on my A3000 (with OS1.3 and OS2.04), and it works on OS3.x, too. I don’t know about OS1.2 (does it differ significantly from 1.3?), but give it a try!

Memory requirements: the program itself doesn’t need much (some Chip memory for the mouse pointer and graphics and other memory — which can also be Chip memory — for the actual code), about a total of 150 kilobytes. The final memory usage depends on you: whether you want to listen to the music or not!

Lines can be installed on harddrive and does not need any additional setting up — just copy the files (especially ‘Lines’ and ‘Lines.info’ or NTSC equivalents) to the destination. Lines will use the same directory when reading or writing settings and hiscores (it locks the directory at startup; if the lock fails, then ‘S:’ is tried as an alternative).

5 First steps

Lines can be started from CLI or Workbench. From CLI, you can give it as an argument the name of the MED module you want to use. If it appears to be packed with PowerPacker, it is unpacked to your Ram Disk and loaded from there. From Workbench, you have two ways to select the music: the first is to give it as a tooltype SONG (i.e. set tooltype ‘SONG=device:path/modulename’ using Info or Information from the Workbench menus), the second is to click first on the icon of the module and then, pressing Shift, to start Lines double-clicking on its icon.

The original archive doesn’t include any musics, because I’m not so good to compose (it is allowed to distribute the archive with some MED modules provided that they are public domain or freely distributable).

Lines is completely played with the mouse, so there is no need for a joystick.

At startup, Lines tries to read its settings and hiscore files (named ‘Lines.Settings’ and ‘Lines.TopLiners’). If settings are not found, then the default ones are used. If environment variable USER exists, its contents are used as a default name, otherwise a hardcoded name is set (it can be changed from the program, of course).

After loading you will see the game screen and logo at the centre. Click on MORE gadget. You will see an info page which shortly tells you the rules and the idea of Lines. Click on OKAY gadget to continue.

Now you are ready to play but let’s first study all the menus.

6 Menus

Click on the right mouse button and move pointer to window title to see the menus. There are three main menus named Project, Edit and Settings. In the Project menu you will find items

New Game, Top Liners, Print, About and Quit, in the Edit menu items Help and Undo and in the

Settings menu items Setup and Music. Almost every item can be used by its first letter (e.g. About by pressing ’A’), either upper or lower case. The only exception is item Help, its abbreviation is the HELP-key.

New Game Starts a new game.

Top Liners

Shows the list of the 10 best players.

Print Prints the screen to paper using your current Preferences settings. During printing the screen colours are changed to white and black. You can cancel printing with ESC-key at any time.

About Shows the starting logo and info.

Quit Quits the game (if TopLiners file has been changed, it is saved).

Help Shows a point belonging to a line (normally, it’s at the either end of that line).

Undo Undoes the last selection (see below for more info).

Setup is used to affect Lines’ colours. Click on the desired colour gadget and use sliders to change the colour. You can also set a default name which is automatically placed on Top Liners list. Click on SAVE to save your settings, DEFAULT to choose original (doesn’t restore the name) or USE to use changes without saving.

Music Stops and continues the music (if music was loaded). It is disabled if there is no music.

On the right-hand side of the screen there is a table from where you can see your amount of Lines, Extras, and the Best result.

For quicker use, UNDO, NEW GAME, and MUSIC can be chosen from gadgets below the score table, the functions are the same their menu equivalents have.

7 The rules

And now to the main thing - the idea and rules of Lines.

The idea is simple: draw as many lines as you can!

Of course, there have to be several rules:

RULE ONE:

Only eight directions (horizontal, vertical, and diagonal) are allowed.

RULE TWO:

A line is exactly five points long.

RULE THREE:

To draw a line, there must exist at least four old points which are not used in that direction. The fifth point may be either at the either end of the line or somewhere between the older points. Of course, it is allowed to draw a line through five old, unused points.

RULE THREE, supplement:

If the line is drawn without drawing any new point, you will get a bonus point, called Extras (and a normal point, of course). Later you can set that Extras point in any position you want and perhaps play a little longer.

8 The playing

And how to play? No problem, just click on the point from where you want to start your line (if it is not valid, the screen will flash). Move mouse to the direction you want to draw the line (if it is a possible direction a blue line is shown). By moving mouse around, you can see all the possible lines from that point. When you have made your selection, click again and you have got a line as well as one point.

After setting the starting point, you can cancel the selection by pressing UNDO (you can notice this easily: the gadget is normally disabled and comes enabled only when it can be used. The same is done to the menu item).

To use an Extras, click twice on the same place and the point is plotted. If the position is illegal, screen will flash after the first click (because the program cannot know whether you are really going to use an Extras or just clicking on a wrong position), but after the second click the Extras is set. You can’t undo any Extras!

9 At the end

After a few games you will learn the play tactics and get better results, and you have a new obsession…

LINES!

10 Brief history

The development of Lines started about two years ago and got finished in few months. It was my first Amiga program, and mainly done to learn C language (I had just got my registered DICE). As time passed, some bugs arose and were fixed. Until now, Lines has been just for my private use. During the winter vacation I had time to ‘polish’ the game (although software is never final…) and decided to release it to the public.

11 The author

Send fees, suggestions and bug reports (none known :–) to the following address:

     Mika Kortelainen
     Petäjätie 13
     FIN-35100 Orivesi as.

I can also be reached through email at the following addresses:

     k125374@cc.tut.fi

and

     k125374@korppi.cs.tut.fi

12 The future

Some enhancements I have in my mind:

1. the game in a Workbench window

2. super bitmap window (i.e. scrollable game board)

3. load/save game

4. use of Screen database (or cloning of Workbench screen), so that users of various graphics boards could get some benefit

5. time game (a certain time period in which a line is to be drawn; otherwise, one line (or point) is subtracted)